OpenStudy (anonymous):

Solve , 5cosx-3sinx=5 .

OpenStudy (anonymous):

\[\frac{\sqrt{5^2 + 3^2} \left(5 \cos{x} - 3\sin{x}\right)}{\sqrt{5^2 + 3^2}}=5\] \[\cos \alpha = \frac{5}{\sqrt{5^2 + 3^2}}, \sin \alpha = \frac{3}{\sqrt{5^2 + 3^2}}\] \[\cos \alpha \cos x - \sin \alpha \sin x = \frac{5}{\sqrt{5^2 + 3^2}} \] \[cos \left( \alpha + x\right) = \frac{5}{\sqrt{5^2+3^2}}\]

OpenStudy (anonymous):

\[x = \cos^{-1} \frac{5}{5^2 + 3^2} - \alpha\]